OpenStudy (anonymous):

you have 5/6 cup of cocoa in your cupboard .A recipe for cookies calls for 2/5 cup of cocoa .How much cocoa would you have left if you made the cookies ?

OpenStudy (anonymous):

(5/6)-(2/5)=x (25/30)-(12/30)=(13/30) 13/30 cannot simplify